OpenStudy (anonymous):

Determine which ordered pairs are solutions to the inequality, y < 2x - 8. A (5, -10) B (3, -6) C (-4, 0) D (10, 5)

OpenStudy (anonymous):

@BunnyBree

OpenStudy (anonymous):

@andrewhaze

OpenStudy (anonymous):

A. (5, -10) y < 2x - 8. -10 < 2(5) - 8 -10 < 10 - 8 -10 < 2 <--- This is true. B. (3, -6) y < 2x - 8. -6 < 2(3) - 8 -6 < 6 - 8 -6 < -2 <-- This is also true. C. (-4, 0) y < 2x - 8. 0 < 2(-4) - 8 0 < -8 - 8 0 < -16 <--- Not true. D. (10, 5) y < 2x - 8. 5 < 2(10) - 8 5 < 20 - 8 5 < 12 <--- This is true. The only ordered pair that doesn't work is C. (-4,0)

OpenStudy (anonymous):

thanks thanks thanks

OpenStudy (anonymous):

i have 2 more questions can you help

OpenStudy (anonymous):

Sure. :D

OpenStudy (anonymous):

Determine the x-intercept and y-intercept of 3x - 6y = 12. A (9, 0) and (0, -3) B (4, 0) and (0, -2) C (-4, 0) and (0, 4)

OpenStudy (andrewhaze):

im not good at these

OpenStudy (anonymous):

3x- 6y =12 The X-intercept occurs where Y=0, so: 3x- 6(0) =12 3x = 12 x = 4 So, the X intercept is (4,0). The Y-intercept occurs where X=0. so: 3(0)-6y=12 -6y =12 y=-2 So the Y intercept is (0,-2) Hence B is the answer.

OpenStudy (anonymous):

OK thanks last question

OpenStudy (anonymous):

Which inequality has the following graph?